# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ B.1 Proof of Lemma 1
The first inequality follows simply from $\Pi^{\mathrm{C}}_{\mathrm{H}}\supset\Pi^{\mathrm{C}}_{\mathrm{M}}\supset\Pi^{\mathrm{C}}_{\mathrm{S}}$. For the second inequality, recall the inclusion relation (2.10). Thus, for any $\pi\in\Pi=\Pi^{\mathrm{C}}_{\mathrm{S}},\Pi^{\mathrm{C}}_{\mathrm{M}},\Pi^{\mathrm{C}}_{\mathrm{H}}$
This implies the second inequality. ∎
B.2 Proof of Proposition 1
Proof.
We consider the distributionally robust Bellman operator $\mathcal{B}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S}$ defined as
$\mathcal{B}(u)(s):=\sup_{d\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u],\quad\forall s\in S.$
$\left\|\mathcal{B}(u)(s)-\mathcal{B}(v)(s)\right\|_{\infty}$ $\stackrel{{\scriptstyle(i)}}{{\leq}}\max_{s\in S}\sup_{d\in\mathcal{Q}}\left|\left(\inf_{p_{s}\in\mathcal{P}_{s}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u]\right)-\left(\inf_{p_{s}\in\mathcal{P}_{s}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[v]\right)\right|$
$=\gamma\max_{s\in S}\sup_{d\in\mathcal{Q}}\left|\inf_{p_{s}\in\mathcal{P}_{s}}d\otimes p_{s}[u]+\sup_{p_{s}\in\mathcal{P}_{s}}-d\otimes p_{s}[v]\right|$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\gamma\max_{s\in S}\sup_{d\in\mathcal{Q}}\max\left\{{\left|\sup_{p_{s}\in\mathcal{P}_{s}}d\otimes p_{s}[u-v]\right|,\left|\inf_{p_{s}\in\mathcal{P}_{s}}d\otimes p_{s}[u-v]\right|}\right\}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\gamma\left\|u-v\right\|_{\infty}$
where $(i)$ follows from the inequality $|\sup_{x}f(x)-\sup_{x}g(x)|\leq\sup_{x}|f(x)-g(x)|$, $(ii)$ is due to $\inf_{x}[f+g]\leq\inf_{x}f+\sup_{x}g\leq\sup_{x}[f+g]$, and $(iii)$ is because $d\otimes p_{s}$ is a probability measure. Therefore, by the Banach fixed point theorem, there exists unique $u^{*}\in\mathbb{R}^{S}$ s.t. $\mathcal{B}(u^{*})=u^{*}$. Moreover,
B.3 Proof of Proposition 2
Proof.
As explained before, we choose $T$ large so that for all $\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}$ and $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$, $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$. Fix an arbitrary $\pi=(d_{1},d_{2},\dots)\in\Pi^{\mathrm{C}}_{\mathrm{H}}$. Recall that by definition of $\Pi^{\mathrm{C}}_{\mathrm{H}}$, for fixed $h_{k}\in\mathbf{H}_{k}$, $d_{k}(h_{k})\in\mathcal{Q}$, where $\mathcal{Q}$ is assumed to be convex. We construct the sequence of Markovian adversaries $\left\{{\kappa^{(-1)},\kappa^{(0)},\dots,\kappa^{(T)}}\right\}$ as follow. Fix any homogeneous adversary $\kappa^{\prime}$ and let $\kappa^{(-1)}=\kappa^{\prime}=(\kappa^{\prime},\kappa^{\prime},\dots)$. Then recursively construct
for some Markovian $\psi_{k+1}(g_{k+1})=\psi_{k+1}(s_{k+1},a_{k+1})$ that we will specify later. This leads to
for all $k=0,\dots,T$.
Now we specify the recursive procedure to construct $\left\{{\kappa^{(k)}}\right\}$. Note that $\kappa^{(-1)}$ is fully specified. Fix any $0\leq k\leq T$, suppose that we have specified $\kappa^{(k-1)}$ (or equivalently, $\psi_{0},\dots,\psi_{k-1}$). Let us define the probability measures $\left\{{\Delta_{k}(s)\in\mathcal{P}(\mathcal{A}):s\in S}\right\}$ by
$\Delta_{k-1}(s)[a]:=E_{\mu}^{\pi,\kappa^{(k-1)}}\left[d_{k-1}(H_{k-1})[a]\middle|X_{k-1}=s\right].$ (B.3)
Since $d_{k-1}(h_{k-1})\in\mathcal{Q}$ where $\mathcal{Q}$ is convex,
$\Delta_{k-1}(s)[\cdot]$ $=\sum_{h_{k-1}\in\mathbf{H}_{k-1}}E_{\mu}^{\pi,\kappa^{(k-1)}}\left[d_{k-1}(H_{k-1})[\cdot]\mathds{1}\left\{{H_{k-1}=h_{k-1}}\right\}\middle|X_{k-1}=s\right]$
$=\sum_{g_{k-2}\in\mathbf{G}_{k-2}}d_{k-1}(g_{k-2},s)[\cdot]E_{\mu}^{\pi,\kappa^{(k-1)}}\left[\mathds{1}\left\{{G_{k-2}=g_{k-2}}\right\}\middle|X_{k-1}=s\right].$
which is a convex combination of $d_{k-1}(g_{k-2},s)\in\mathcal{Q}$. Therefore, $\Delta_{k-1}(s)\in\mathcal{Q}$
Fix any $\delta>0$. Since for all $s\in S$
$u^{*}(s)$ $=\sup_{d\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u^{*}]$ (B.4)
$\geq\inf_{p_{s}\in\mathcal{P}_{s}}\Delta_{k-1}(s)[r(s,\cdot)]+\gamma\Delta_{k-1}(s)\otimes p_{s}[u^{*}]$
there exists $\left\{{q_{s}\in\mathcal{P}_{s}:s\in S}\right\}$ s.t.
$u^{*}(s)+\delta\geq\Delta_{k-1}(s)[r(s,\cdot)]+\gamma\Delta_{k-1}(s)\otimes q_{s}[u^{*}].$
Then, for each $s$ and $q_{s}=\left\{{q_{s,a}:a\in A}\right\}$ as above, we define $\psi_{k-1}(s,a)=\left\{{q_{s,a}:s,a\in S\times A}\right\}$. Then
$\Delta_{k-1}(s)[r(s,\cdot)]+\gamma\Delta_{k-1}(s)\otimes\psi_{k-1}(s,\cdot)[u^{*}]\leq u^{*}(s)+\delta.$ (B.5)
Doing this for $k=1,\dots,T$ will complete the construction.
Now, we claim and defer the proof of the inequality
$v_{\mu}^{\pi}(k,\kappa^{(k)})\leq v_{\mu}^{\pi}(k-1,\kappa^{(k-1)})+\gamma^{k-1}\delta$ (B.6)
for all $1\leq k\leq T$. With (B.6), we see that
$v_{\mu}^{\pi}(T,\kappa^{(T)})$ $\leq v_{\mu}^{\pi}(T-1,\kappa^{(T-1)})+\gamma^{T-1}\delta$
$\leq v_{\mu}^{\pi}(0,\kappa^{(0)})+\sum_{k=0}^{T-1}\gamma^{k}\delta$
$\leq v_{\mu}^{\pi}(0,\kappa^{(0)})+\frac{\delta}{1-\gamma}$
$\leq v_{\mu}^{\pi}(T,\kappa^{(T)})+\frac{\epsilon}{2}$
$\leq\mu[u^{*}]+\epsilon.$
Finally, it remains to prove the claim (B.6). We begin with analyzing the following quantity
$\quad E_{\mu}^{\pi,\kappa^{(k)}}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})\right]$ (B.7)
$=E_{\mu}^{\pi,\kappa^{(k)}}\left[d_{k-1}(H_{k-1})[r(X_{k-1},\cdot)]+\gamma d_{k-1}(H_{k-1})\otimes\psi_{k-1}(X_{k-1},\cdot)[u^{*}]\right]$
Notice that
$\quad E_{\mu}^{\pi,\kappa^{(k)}}\left[d_{k-1}(H_{k-1})\otimes\psi_{k-1}(X_{k-1},\cdot)[u^{*}]\middle|X_{k-1}=s\right]$
$=E_{\mu}^{\pi,\kappa^{(k)}}\left[\sum_{s^{\prime}\in S}\sum_{a\in A}d_{k-1}(H_{k-1})[a]\psi_{k-1}(s,a)(s^{\prime})u^{*}(s^{\prime})\middle|X_{k-1}=s\right]$
$=\sum_{s^{\prime}\in S}\sum_{a\in A}\psi_{k-1}(s,a)[s^{\prime}]u^{*}(s^{\prime})E_{\mu}^{\pi,\kappa^{(k)}}\left[d_{k-1}(H_{k-1})[a]\middle|X_{k-1}=s\right]$
=$\stackrel{{\scriptstyle(i)}}{{=}}\sum_{s^{\prime}\in S}\sum_{a\in A}\psi_{k-1}(s,a)[s^{\prime}]u^{*}(s^{\prime})E_{\mu}^{\pi,\kappa^{(k-1)}}\left[d_{k-1}(H_{k-1})[a]\middle|X_{k-1}=s\right]$
$=\sum_{s^{\prime}\in S}\sum_{a\in A}\psi_{k-1}(s,a)[s^{\prime}]u^{*}(s^{\prime})\Delta_{k-1}(s)[a]$
$=\Delta_{k-1}(s)\otimes\psi_{k-1}(s,\cdot)[u^{*}]$
where $(i)$ follows from the observation that the law of $H_{k-1}$ is not dependent on the adversarial decision rules after time $k-1$. Apply this to (B.7), we have that
$\quad E_{\mu}^{\pi,\kappa^{(k)}}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})|X_{k-1}\right]$ (B.8)
$=\Delta_{k-1}(X_{k-1})[r(X_{k-1},\cdot)]+\gamma\Delta_{k-1}(X_{k-1})\otimes\psi_{k-1}(X_{k-1},\cdot)[u^{*}]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}u^{*}(X_{k-1})+\delta.$
where $(i)$ follows from (B.5). With this result, we can bound
$v_{\mu}^{\pi}(k,\kappa^{(k)})$ $=E_{\mu}^{\pi,\kappa^{(k)}}\left[\sum_{j=0}^{k-2}\gamma^{j}r(X_{j},A_{j})\right]+\gamma^{k-1}E_{\mu}^{\pi,\kappa^{(k)}}E_{\mu}^{\pi,\kappa^{(k)}}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})\middle|X_{k-1}\right]$
$\leq E_{\mu}^{\pi,\kappa^{(k)}}\left[\sum_{j=0}^{k-2}\gamma^{j}r(X_{j},A_{j})\right]+\gamma^{k-1}E_{\mu}^{\pi,\kappa^{(k)}}u^{*}(X_{k-1})+\gamma^{k-1}\delta$
=$\stackrel{{\scriptstyle(i)}}{{=}}E_{\mu}^{\pi,\kappa^{(k-1)}}\left[\sum_{j=0}^{k-2}\gamma^{j}r(X_{j},A_{j})+\gamma^{k-1}u^{*}(X_{k-1})\right]+\gamma^{k-1}\delta$
$=v_{\mu}^{\pi}(k-1,\kappa^{(k-1)})+\gamma^{k-1}\delta$
as claimed, where $(i)$ follows from the quantity inside the expectation is a bounded function of $H_{k-1}$. ∎
B.4 Proof of Proposition 3
Proof.
It is equivalent to show that for any $\pi\in\Pi^{\mathrm{C}}_{\mathrm{U}}$
$\inf_{\kappa\in\mathrm{K}^{\mathrm{C}}_{\mathrm{U}}}v(\mu,\pi,\kappa)\leq\mu[u^{*}];$
for $\mathrm{U}=\mathrm{H},\mathrm{M},\mathrm{S}$, respectively.
Fix any $\epsilon>0$. Given a policy $\pi=(d_{0},d_{1},\dots)\in\Pi^{\mathrm{C}}_{\mathrm{H}}$ (or $\Pi^{\mathrm{C}}_{\mathrm{M}}$, $\Pi^{\mathrm{C}}_{\mathrm{S}}$), we prove this by constructing an adversarial policy $\kappa=(\kappa_{0},\kappa_{1},\dots)\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ (or $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ resp.), so that $v(\mu,\pi,\kappa)\leq\mu[u^{*}]+\epsilon.$
For all $k\geq 0$ and $h_{k}=(g_{k-1},s)\in\mathbf{H}_{k}$, as $u^{*}$ solves (3.1),
$d_{k}(h_{k})[r(s,\cdot)]+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}d_{k}(h_{k})\otimes p_{s}[u^{*}]\leq u^{*}(s).$
For any $\delta>0$, we can always choose $\kappa_{k}(h_{k},\cdot)\in\mathcal{P}_{s}$ so that
$d_{k}(h_{k})[r(s,\cdot)]+\gamma d_{k}(h_{k})\otimes\kappa_{k}(h_{k},\cdot)[u^{*}]\leq u^{*}(s)+\delta$ (B.9)
for all $k\geq 0$ and $h_{k}=(g_{k-1},s)\in\mathbf{H}_{k}$. Moreover, if $d_{k}(h_{k})=d_{k}(s)$ is Markov (or $d_{k}(h_{k})=d(s)$ is Markov time-homogeneous), we can always choose $\kappa_{k}(h_{k},\cdot)=\kappa_{k}(s,\cdot)$ (or $\kappa_{k}(h_{k},\cdot)=\kappa(s,\cdot)$ Markov time-homogeneous) so that (B.9) holds for all $k\geq 0$ and $h_{k}=(g_{k-1},s)\in\mathbf{H}_{k}$. Define $\kappa=(\kappa_{0},\kappa_{1},\dots)\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ (or $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ resp.) with $\left\{{\kappa_{k}:k\geq 0}\right\}$ specified above.
Next, we consider
@@ B.4 Proof of Proposition 3
$=E_{\mu}^{\pi,\kappa}d_{k-1}(H_{k-1})[r(X_{k-1},\cdot)]+\gamma d_{k-1}(H_{k-1})\otimes\kappa_{k-1}(H_{k-1},\cdot)[u^{*}]$
$\leq E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})+\delta.$
With this bound, we choose $T$ large so that for all $\pi,\kappa$, $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$. Then,
$=v_{\mu}^{\pi}(T-1,\kappa)+\gamma^{T-1}\delta$
[…]
$\leq v_{\mu}^{\pi}(0,\kappa)+\sum_{k=0}^{T-1}\gamma^{k}\delta$
$\leq\mu[u^{*}]+\frac{\delta}{1-\gamma}.$
$v^{*}(\mu,\pi,\kappa)\leq v^{\pi}_{\mu}(T,\kappa^{\prime})+\frac{\epsilon}{2}\leq\mu[u^{*}]+\epsilon.$
Recall that $\epsilon>0$ and $\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}$ (or $\Pi^{\mathrm{C}}_{\mathrm{M}}$, $\Pi^{\mathrm{C}}_{\mathrm{S}}$) are arbitrary, with corresponding $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ (or $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ resp.) This completes the proof of the proposition. ∎
B.5 Proof of Proposition 4
Proof.
Since $u^{*}$ is the solution to (3.1), for any fixed $\delta>0$, there exists Markov time-homogeneous decision rule $\Delta:S\rightarrow\mathcal{Q}$ s.t.
$u^{*}(s)\leq\Delta(s)[r(s,\cdot)]+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}\Delta(s)\otimes p_{s}[u^{*}]+\delta.$ (B.10)
Let $\pi=(\Delta,\Delta,\dots)\in\Pi^{\mathrm{C}}_{\mathrm{S}}$. We consider for any $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$\gamma E_{\mu}^{\pi,\kappa}\left[u^{*}(X_{k})\right]$ $=\gamma E_{\mu}^{\pi,\kappa}E_{\mu}^{\pi,\kappa}\left[u^{*}(X_{k})\middle|G_{k-1}\right]$
$=\gamma E_{\mu}^{\pi,\kappa}\kappa_{k-1}(G_{k-1})[u^{*}]$
$=\gamma E_{\mu}^{\pi,\kappa}E_{\mu}^{\pi,\kappa}\left[\kappa_{k-1}(G_{k-1})[u^{*}]\middle|H_{k-1}\right]$
$=\gamma E_{\mu}^{\pi,\kappa}\sum_{a\in A}\Delta(X_{k-1})[a]\kappa_{k-1}(H_{k-1},a)[u^{*}]$
Now, if we define the random adversarial decision rule $K_{k-1,X_{k-1}}:=\left\{{\kappa_{k-1}(H_{k-1},a):a\in A}\right\}$. Then, as $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$, we have $K_{k-1,X_{k-1}}\in\mathcal{P}_{X_{k-1}}$ w.p.1. Moreover, by the notation defined in (2.14),
$\gamma E_{\mu}^{\pi,\kappa}\left[u^{*}(X_{k})\right]$ $=\gamma E_{\mu}^{\pi,\kappa}\Delta(X_{k-1})\otimes K_{k-1,X_{k-1}}[u^{*}]$
$\geq E_{\mu}^{\pi,\kappa}\gamma\inf_{p_{X_{k-1}}\in\mathcal{P}_{X_{k-1}}}\Delta(X_{k-1})\otimes p_{X_{k-1}}[u^{*}]$
$\stackrel{{\scriptstyle(i)}}{{\geq}}E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})-\Delta(X_{k-1})[r(X_{k-1},\cdot)]-\delta$
=$\stackrel{{\scriptstyle(ii)}}{{=}}E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})-E_{\mu}^{\pi,\kappa}r(X_{k-1},A_{k-1})-\delta$
where $(i)$ follows from (B.10) and $(ii)$ can be seen by first condition on $G_{k-2}$ and noting that $\Delta$ is history independent given $X_{k-1}$.
Fix any $\epsilon>0$, let $\pi$ be defined as above. Recall the definition of $v_{\mu}^{\pi}(T,\kappa)$ in (B.2). We still choose $T$ large so that for all $\kappa$, $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$. Then, for any $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$\geq E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})\right]+\gamma^{T-1}\left[E_{\mu}^{\pi,\kappa}u^{*}(X_{T-1})-E_{\mu}^{\pi,\kappa}r(X_{T-1},A_{T-1})-\delta\right]$
$=v_{\mu}^{\pi}(T-1,\kappa)-\gamma^{T-1}\delta$
[…]
$\geq v_{\mu}^{\pi}(0,\kappa)-\sum_{k=0}^{T-1}\gamma^{k}\delta$
$\geq\mu[u^{*}]-\frac{\delta}{1-\gamma}.$
By definition of $\inf$ and that $\pi\in\Pi^{\mathrm{C}}_{\mathrm{S}}$, there exists $\kappa^{\prime}$ s.t.
$v(\mu,\pi,\kappa^{\prime})-\frac{\epsilon}{4}\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}}v(\mu,\pi,\kappa)\leq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}).$
Then,
00 0 $\stackrel{{\scriptstyle(i)}}{{\leq}}\mu[u^{*}]-v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ (B.11)
$\leq\mu[u^{*}]-v(\mu,\pi,\kappa^{\prime})+\frac{\epsilon}{4}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\mu[u^{*}]-v^{\pi}_{\mu}(T,\kappa^{\prime})+\frac{3\epsilon}{4}$
$\leq\frac{\delta}{1-\gamma}+\frac{3\epsilon}{4}.$
where $(i)$ follows from Proposition 3, and $(ii)$ is due to the choice of $T$. Since $\delta$ can be chosen as $(1-\gamma)\epsilon/4$, this implies $\mu[u^{*}]=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ as $\epsilon>0$ is arbitrary. ∎
B.6 Proof of Propositions 5 and 6
Proof of Proposition 5.
Since $u^{*}$ satisfies (3.2), for any $\delta>0$, there exists adversarial decision rule $\psi=\left\{{\psi(s,\cdot)\in\mathcal{P}_{s}:s\in S}\right\}$ s.t.
$u^{*}(s)\geq\sup_{d\in\mathcal{Q}}d[r(s,\cdot)]+\gamma d\otimes\psi(s,\cdot)[u^{*}]-\delta.$
Let $\kappa=(\psi,\psi,\dots)\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\pi=(d_{0},d_{1},d_{2}\dots)\in\Pi^{\mathrm{C}}_{\mathrm{H}}$ be any controller policy. Consider
$=E_{\mu}^{\pi,\kappa}d_{k-1}(H_{k-1})[r(X_{k-1},\cdot)]+\gamma d_{k-1}(H_{k-1})\otimes\psi(X_{k-1},\cdot)[u^{*}]$
$\leq E_{\mu}^{\pi,\kappa}\sup_{d\in\mathcal{Q}}d[r(X_{k-1},\cdot)]+\gamma d\otimes\psi(X_{k-1},\cdot)[u^{*}]$
$\leq E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})+\delta$
@@ B.6 Proof of Propositions 5 and 6
$\leq\mu[u^{*}]+\frac{\delta}{1-\gamma}.$
@@
$\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}}v^{*}(\mu,\pi,\kappa)\leq v^{*}(\mu,\pi,\kappa)\leq v^{\pi}_{\mu}(T,\kappa^{\prime})+\frac{\epsilon}{2}\leq\mu[u^{*}]+\epsilon.$
Since $\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}$ and $\epsilon>0$ are arbitrary, this implies the proposition. ∎
@@
Proof of Proposition 6.
For any $\delta>0$, let $\kappa=(\psi,\psi,\dots)\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ be constructed as above. Then the same proof as above implies that for any $\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}$,
$v_{\mu}^{\pi}(T,\kappa)\leq\mu[u^{*}]+\frac{\delta}{1-\gamma}.$
Therefore, by the choice of $T$ and the definition of sup, there exists $\pi^{\prime}\in\Pi^{\mathrm{C}}_{\mathrm{H}}$ s.t.
$\sup_{\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}}v(\mu,\pi,\kappa)$ $\leq v(\mu,\pi^{\prime},\kappa)+\frac{\epsilon}{4}$
@@
$\leq\mu[u^{*}]+\frac{\delta}{1-\gamma}+\frac{3\epsilon}{4}.$

### v4

@@ B.3.1 Proof of Propositions 4 and 5
Proof of Proposition 4
To prove Proposition 4, we use the min-max equation (3.2) to construct a time-homogeneous adversary $\kappa=\left\{{\psi,\psi,\dots}\right\}$ that upper bounds the control value.
Since $u^{*}$ satisfies (3.2), for any $\delta>0$, there exists a Markov time-homogeneous adversarial decision rule $\psi=\left\{{\psi(\cdot|s,\cdot)\in\mathcal{P}_{s}:s\in S}\right\}$ s.t.
$u^{*}(s)\geq\sup_{\phi\in\mathcal{Q}}E_{\phi,\psi(\cdot|s,\cdot)}[r(s,A_{0})+\gamma u^{*}(X_{1})]-\delta.$
Let $\kappa=(\psi,\psi,\dots)\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\pi=(\pi_{0},\pi_{1},\pi_{2}\dots)\in\Pi_{\mathrm{H}}$ be any controller policy. Consider
@@
$\stackrel{{\scriptstyle(i)}}{{=}}E_{\mu}^{\pi,\kappa}\sum_{a\in A}\sum_{s^{\prime}\in S}\pi_{k-1}(a|H_{k-1})\psi(s^{\prime}|X_{k-1},a)[r(X_{k-1},a)+\gamma u^{*}(s^{\prime})]$
$\leq E_{\mu}^{\pi,\kappa}\sup_{\phi\in\mathcal{Q}}\sum_{a\in A}\sum_{s^{\prime}\in S}\phi(a)\psi(s^{\prime}|X_{k-1},a)[r(X_{k-1},a)+\gamma u^{*}(s^{\prime})]$
$=E_{\mu}^{\pi,\kappa}\left(\sup_{\phi\in\mathcal{Q}}E_{\phi,\psi(\cdot|s,\cdot)}[r(s,A_{0})+\gamma u^{*}(X_{1})]\right)_{s=X_{k-1}}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})+\delta$
where $(i)$ follows from the construction that under $E_{\mu}^{\pi,\kappa}[\cdot|H_{k-1}]$, $A_{k-1},X_{k}\sim E_{\pi_{k}(\cdot|X_{H_{k-1}}),\psi(\cdot|X_{k-1},\cdot)}$, and $(ii)$ follows from the definition of $\psi$.
@@
$\leq E_{\mu}[u^{*}(X_{0})]+\frac{\delta}{1-\gamma}.$
@@
$\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}}v(\mu,\pi,\kappa)\leq v(\mu,\pi,\kappa)\leq v^{\pi}_{\mu}(T,\kappa^{\prime})+\frac{\epsilon}{2}\leq E_{\mu}[u^{*}(X_{0})]+\epsilon.$
Since $\pi\in\Pi_{\mathrm{H}}$ and $\epsilon>0$ are arbitrary, this implies the proposition. ∎
@@
Proof.
Proof of Proposition 5 For any $\delta>0$, let $\kappa=(\psi,\psi,\dots)\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ be constructed as above. Then the same proof as above implies that for any $\pi\in\Pi_{\mathrm{H}}$,
$v_{\mu}^{\pi}(T,\kappa)\leq E_{\mu}[u^{*}(X_{0})]+\frac{\delta}{1-\gamma}.$
Therefore, by the choice of $T$ and the definition of sup, there exists $\pi^{\prime}\in\Pi_{\mathrm{H}}$ s.t.
$\sup_{\pi\in\Pi_{\mathrm{H}}}v(\mu,\pi,\kappa)$ $\leq v(\mu,\pi^{\prime},\kappa)+\frac{\epsilon}{4}$
@@
$\leq E_{\mu}[u^{*}(X_{0})]+\frac{\delta}{1-\gamma}+\frac{3\epsilon}{4}.$